Difffusion Morning Report 05/10/04

ted berg tberg at snf.stanford.edu
Mon May 10 07:44:29 PDT 2004


Good Morning All,
    Looking pretty good this morning. Tylan 1 power supply replaced and 
looks like it is OK ,testing.  Nitride is back up after pull and clean.
LTO is down for scheduled pull and clean.  Burn box cleaned this 
morning.  TEOS 2 is testing-working out minor bugs.  Epi has some 
issues-investigating
Everything else looks OK.ted

Diffusion 	Date: 	5/10/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	yellow 	power supply replaced needs test 	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	MFC is installed testing-trap installed 	 
Tylan 8-SiGe 	Up 	  	6.6/30
Tylan 9-Poly 	Up 	  	27.7/30
Tylan 10-Nitride 	Up 	Particles? 	4.8/30
Tylan 11-BPSG-LTO 	down 	Pull and clean 	0/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	11.2/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	Cleaned this morning 	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 540 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040510/7ca6bc2c/attachment.html>


More information about the diffusion mailing list